OpenStudy (anonymous):

Form a polynomial function with real coefficients having the given: Degree: 4; zeros; -1, 2 and 1-2i

OpenStudy (anonymous):

I just forgot how to set this one up with the 'i' in there I think I can figure it out from there if someone shows me how to set it up

OpenStudy (campbell_st):

well I think you may have missed a zero perhaps is \[1 \pm 2i\]

OpenStudy (dumbcow):

start with \[x = 1-2i\] then work backwards to get it to equal 0 (with no i) \[x-1 = -2i\] \[(x-1)^{2} = (-2i)^{2}\] ...

OpenStudy (campbell_st):

so it can be P(a) = (a +1)(a - 2)(a + (1-2i))(a - (1 -2i))

OpenStudy (anonymous):

thank you !